OpenStudy (anonymous):

d^75 / dx^75 of (sinx)?

OpenStudy (turingtest):

after how many (i.e for which n) derivatives does \[{{d^n}\over{dx^n}}\sin x\]repeat?

OpenStudy (anonymous):

75

OpenStudy (dumbcow):

it repeats every multiple of 5

OpenStudy (anonymous):

It looks like that

OpenStudy (turingtest):

yeah, and u=sinx, u'=cosx, u''=-sinx, u'''=-cosx, u^(4)=sinx so they repeat

OpenStudy (anonymous):

So u^5 = cosx?

OpenStudy (turingtest):

yep